Five hundred children participated in a field demonstration.Their heights averaged 110 cm with a standard deviation of 6 cm.How many children , picked at random,has a height greater than 116 cm?


1
Expert's answer
2021-06-29T03:32:23-0400

"N=500 \\

\mu=110 \\

\sigma=6 \\

P(X>116) = 1 -P(X<116) \\

= 1 - P(Z< \frac{116-110}{6}) \\

= 1- P(Z<1) \\

= 1 -0.3413 \\

= 0.8413"

Number of children, which heights greater than 116 cm =

"= 0.8413 \times 500 \\

= 421"